OpenStudy (anonymous):

A sample space S consists of five simple events with these probabilities: P(E1) = P(E2) = 0.15 P(E3) = 0.4 P(E4) = 2P(E5) a) Find the probabilities for simple events E4 and E5. How do I do this?

jimthompson5910 (jim_thompson5910):

let x = P(E5) so P(E4) = 2*P(E5) = 2*x

jimthompson5910 (jim_thompson5910):

for this distribution to be a true probability distribution, all of the probabilities must add to 1 P(E1) + P(E2) + .... + P(E5) = 1

OpenStudy (anonymous):

right but how do i fugure the probability out from x?

jimthompson5910 (jim_thompson5910):

P(E1) + P(E2) + P(E3) + P(E4) + P(E5) = 1 0.15+0.15+0.4+2x+x = 1 solve for x

OpenStudy (anonymous):

x = 0.10?

jimthompson5910 (jim_thompson5910):

once you find x, you'll know P(E5) double x to get P(E4)

jimthompson5910 (jim_thompson5910):

x = 0.10 is correct

OpenStudy (anonymous):

Thank you Jim :)

jimthompson5910 (jim_thompson5910):

np